A sample of xenon gas at 20°C and o.57o atm is called to a pressure of o.10o atm volume remains constant, what is the final Celsius temperature: atm is called to a pressure of o.1oo atmn. If the If o soo mol of hydrogen gas occupies so.o ml. at asC, what is the pressure in atmospheres: 9. If1.25 mol of oxygen gas exerts a pressure of iz00 mm Hg at a5C, what is the volume liters: 1. A gas occupies 900.0 mL at a temperature of 270 oC. What is the volume at i320°C: What volume would 42.o g of NO. gas occupy at 3.12 atm and 18.0 °C: 2. A gas occupies 12.3 liters at a pressure of 40.0 mm Hg. What is the volume when the pressure is increased to 60.o mm Hg: 3. 4. If 6.50 mol of CO occupy 20.6 L, how many liters will 16.5 mol of CO occupy at the same temperature and pressure:

can I get a step by step explanation on this problem please
Answer #1

o-S구. 343 51-9k o-0S 00mm H 4 C o. 91xo-0921X291-16.9礼 la こ 12. 604. At constant temperature and pressure

V1/n1 = V2/n2

V2 = V1×n2/n1

= 20.6×16.5/6.5

= 52.3 L